# DuetGraph: Coarse-to-Fine Knowledge Graph Reasoning with Dual-Pathway Global-Local Fusion

### v1

@@ Appendix A Proofs of theorems
In this section, we provide the theorem proofs of in method part, including 1) why our proposed dual-pathway global-local fusion model can alleviate the over-smoothing in KG; 2) why our proposed coarse-to-fine reasoning optimization can alleviate the over-smoothing in KG; 3) why our proposed coarse-to-fine reasoning optimization can improve the quality of KG reasoning.
@@
Let entity initial representation is denoted as $X^{(0)}\in\mathbb{R}^{n\times d}$, symmetrically normalized adjacency matrix in message passing networks is denoted as $A\in\mathbb{R}^{n\times n}$. The attention matrix computed by a single layer of global attention is denoted as $P\in\mathbb{R}^{n\times n}$. We construct a weight matrix of one-pathway model stacked with $L$ message passing layers and a transformer layer defined as:
@@
where $f:\mathbb{R}^{d}\to\mathbb{R}$ denotes the MLP and $L_{f}$ is the Lipschitz constant [44]. Since $A$ is a symmetrically normalized adjacency matrix, spectral theory ensures that its eigenvalues $\{\lambda_{i}\}^{n}_{i=1}$ satisfy $11=\lambda_{1}>\lambda_{2}\geq\cdots\geq\lambda_{n}>-$ 1. Hence, we conclude that largest eigenvalue of $A^{L}$ is equal to 1, and its largest singular value $\sigma_{\max}(A^{L})$ is equal to 1.
@@
Therefore, as long as the learnable parameter $\alpha$ is less than $\frac{\sigma_{\max}(\mathcal{M}_{\mathrm{D}})+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}{1+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}$, $\sigma_{\max}(\mathcal{M}_{D})$ will necessarily be greater than $\sigma_{\max}(\mathcal{M}_{O})$. The result of Equation 11 indicates that as the number of iteration $\ell$ increases, the score gap between any two entities will decrease exponentially with respect $\sigma_{\max}(\mathcal{M})$. This implies that If $\alpha<\frac{\sigma_{\max}(\mathcal{M}_{\mathrm{D}})+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}{1+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}$, the score gap upper bound of the the dual-pathway model is greater than that of the one-pathway model and the dual-pathway model shows a slower decrease of the score gap upper bound. Consequently, dual-pathway model effectively mitigates the oversmoothing. ∎
A.2 Coares-to-Fine Reasoning Optimization Alleviates the over-smoothing in KG.
@@
In our implementation, $N_{h}$ is less than 10 and the number of entity in all datasets is more than 10000. Then, we have $\frac{N_{l}}{N_{h}}>1000$, and the function $f(x)=\frac{x}{x^{2}+1}$ is monotonically decreasing for $x>1$. Therefore,
@@
We have demonstrate that, in our coarse-to-fine reasoning optimization, the expected score gap between the high-score and low-score subsets is at least 0.1 times the standard deviation. In comparision, other baseline methods (as shown in Figure 1) exhibit score gaps between correct and incorrect answers are typically less than $0.02\sigma$. This demonstrates that our optimization can amplifying the score gap, thus mitigating over-smoothing.
@@
In coarse-grained reasoning, the candidate entities are divided into two subsets, the high-score subset is denoted as $\mathcal{T}^{high}$. The highest-scoring entity in each subset, as computed by our proposed dual-pathway fusion model, is denoted as:
@@
where the $s(\cdot)$ denotes score computing by dual-pathway fusion model. Let $P_{\Delta}$ denote the probability that the difference between $e_{h}$ and $e_{l}$ is less than or equal to $\Delta$, i.e. $P_{\Delta}$ = $P(e_{l}-e_{h}\leq\Delta)$
@@
Therefore, we can obtain that $P(B\mid A)$geq $P^{\prime}$ which leads to $P\geq P^{\prime}$, thus demonstrating that the probability that coarse-to-fine reasoning optimization accurately infers the correct answer is more than the probability that dual-pathway fusion model without coarse-to-fine optimization correctly infers the correct answer. ∎
@@ Appendix B Time complexity computation
The coarse-to-fine reasoning stage has two additional operations of coarse-grained reasoning and sorting all entities compared to one-stage. In coarse-to-fine reasoning, parallel reasoning with coarse model and fine model. Morever, the time complexity of the coarse model we employ is $\mathcal{O}(|\mathcal{E}|d)$ where $\left|\mathcal{E}\right|$ denote the number of triplets and $d$ is the dimension of entity representation. The time complexity of this sorting process is $\mathcal{O}(|\mathcal{V}|log|\mathcal{V}|)$, where $\left|\mathcal{V}\right|$ denote the number of entities. The one-stage reasoning only includes dual-pathway fusion model. Therefore, the overall time complexity of the coarse-to-fine reasoning stage is $\mathcal{O}(\max(L_{m}(\left|\mathcal{E}\right|d+\left|\mathcal{V}\right|d^{2}),L_{t}\left|\mathcal{V}\right|d^{2})+\left|\mathcal{V}\right|log\left|\mathcal{V}\right|)$.
@@ Appendix C Additional baseline discussion
C.1 DuetGraph v.s. Methods based on pre-trained language models.
We observe that language model-based reasoning methods such as SimKGC [52] and MoCoKGC [57] achieve unusually high results on WN18RR, but perform poorly on other datasets, as shown in Table 5. To better understand this phenomenon, we take these two methods as representative examples for further analysis. We note that WN18RR is derived from WordNet, a large lexical database of English that naturally encodes rich semantic relations between words. Pretrained language models are well-suited to capturing such general semantic information, which may explain their strong performance on WN18RR. In contrast, FB15k-237 involves more domain-specific relational knowledge, which poses greater challenges for these models, leading to weaker performance(as shown in Table 5).
@@ Appendix D Experimental Details
D.1 Dataset Statistics
@@ D.2 Hyperparameters Setup
[…]
@@
We analyze the impact of the hyper-parameter $k$, which denotes the number of entities in the high-score subset. We conduct experiments on the validation sets of all transductive datasets using our model. We set $k$ to {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15}. See Table 9. We ultimately select $k=7$ for WN18RR and $k=4$ for the other datasets, as these settings yield relatively high and stable results across three key metrics—MRR, Hits@1, and Hits@10—rather than optimizing a single metric in isolation. We also conduct the same experiments on 4 inductive datasets (As shown in in Figure 5).
@@
Additionaly, We analyze the impact of the hyper-parameter, the decision threshold $\Delta$. We run experiments on all transductive datasets with our model. We set $\Delta$ to {0, 0.5, 1, 1.5, 2, 2.5, 3, 3.5, 4, 4.5, 5, 5.5, 6, 6.5, 7, 7.5, 8, 8.5, 9, 9.5, 10}. For FB15k-237 and WN18RR, we set $\Delta$ as 8 and set $\Delta$ as 5 for NELL-995 and YAGO3-10.
@@
Appendix E Limitations and Broader Impacts
E.1 Limitations

### v2

@@ Appendix A Proofs of Theorems
In this section, we provide the theorem proofs in method part, including 1) why our proposed dual-pathway global-local fusion model can alleviate the over-smoothing in KG; 2) why our proposed coarse-to-fine reasoning optimization can alleviate the over-smoothing in KG; 3) why our proposed coarse-to-fine reasoning optimization can improve the quality of KG reasoning.
@@
Let entity initial representation be denoted as $X^{(0)}\in\mathbb{R}^{n\times d}$, symmetrically normalized adjacency matrix in message passing networks is denoted as $A\in\mathbb{R}^{n\times n}$. The attention matrix computed by a single layer of global attention is denoted as $P\in\mathbb{R}^{n\times n}$. We construct a weight matrix of one-pathway model stacked with $L$ message passing layers and a transformer layer defined as:
@@
where $f:\mathbb{R}^{d}\to\mathbb{R}$ denotes the MLP and $L_{f}$ is the Lipschitz constant [46]. Since $A$ is a symmetrically normalized adjacency matrix, spectral theory ensures that its eigenvalues $\{\lambda_{i}\}^{n}_{i=1}$ satisfy $1=\lambda_{1}>\lambda_{2}\geq\cdots\geq\lambda_{n}>-1$. Hence, we conclude that largest eigenvalue of $A^{L}$ is equal to 1, and its largest singular value $\sigma_{\max}(A^{L})$ is equal to 1.
@@
Therefore, as long as the learnable parameter $\alpha$ is less than $\frac{\sigma_{\max}(\mathcal{M}_{\mathrm{D}})+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}{1+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}$, $\sigma_{\max}(\mathcal{M}_{D})$ will necessarily be greater than $\sigma_{\max}(\mathcal{M}_{O})$. The result of Equation 11 indicates that as the number of iterations $\ell$ increases, the score gap between any two entities will decrease exponentially with respect $\sigma_{\max}(\mathcal{M})$. This implies that If $\alpha<\frac{\sigma_{\max}(\mathcal{M}_{\mathrm{D}})+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}{1+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}$, the score gap upper bound of the dual-pathway model is greater than that of the one-pathway model and the dual-pathway model shows a slower decrease in the score gap upper bound. Consequently, dual-pathway model effectively mitigates the oversmoothing. ∎
A.2 Coarse-to-Fine Reasoning Optimization Alleviates the over-smoothing in KG.
@@
In our implementation, $N_{h}$ is less than 10 and the number of entities in all datasets is more than 10000. Then, we have $\frac{N_{l}}{N_{h}}>1000$, and the function $f(x)=\frac{x}{x^{2}+1}$ is monotonically decreasing for $x>1$. Therefore,
@@
We have demonstrated that, in our coarse-to-fine reasoning optimization, the expected score gap between the high-score and low-score subsets is at least 0.1 times the standard deviation. In comparision, other baseline methods (as shown in Figure 1) exhibit score gaps between correct and incorrect answers are typically less than $0.02\sigma$. This demonstrates that our optimization can amplify the score gap, thus mitigating over-smoothing.
@@
In coarse-grained reasoning, the candidate entities are divided into two subsets, the high-score subset is denoted as $\mathcal{T}^{high}$. The highest-score entity in each subset, as computed by our proposed dual-pathway fusion model, is denoted as:
@@
where the $s(\cdot)$ denotes score computing by dual-pathway fusion model. Let $P_{\Delta}$ denote the probability that the difference between $e_{h}$ and $e_{l}$ is less than or equal to $\Delta$, i.e. $P_{\Delta}$ = $P(e_{l}-e_{h}\leq\Delta)$.
@@
Therefore, we can obtain that $(B\mid A) \geq PP^{\prime}$ which leads to $P\geq P^{\prime}$, thus demonstrating that the probability that coarse-to-fine reasoning optimization accurately infers the correct answer is more than the probability that dual-pathway fusion model without coarse-to-fine optimization correctly infers the correct answer. ∎
@@ Appendix B Time Complexity Computation
The coarse-to-fine reasoning stage has two additional operations of coarse-grained reasoning and sorting all entities compared to one-stage. In coarse-to-fine reasoning, parallel reasoning with coarse model and fine model. Moreover, the time complexity of the coarse model we employ is $\mathcal{O}(|\mathcal{E}|d)$ where $\left|\mathcal{E}\right|$ denote the number of triplets and $d$ is the dimension of entity representation. The time complexity of this sorting process is $\mathcal{O}(|\mathcal{V}|log|\mathcal{V}|)$, where $\left|\mathcal{V}\right|$ denote the number of entities. The one-stage reasoning only includes dual-pathway fusion model. Therefore, the overall time complexity of the coarse-to-fine reasoning stage is $\mathcal{O}(\max(L_{m}(\left|\mathcal{E}\right|d+\left|\mathcal{V}\right|d^{2}),L_{t}\left|\mathcal{V}\right|d^{2})+\left|\mathcal{V}\right|log\left|\mathcal{V}\right|)$.
@@ Appendix C Additional Baseline Discussion
C.1 DuetGraph vs. Methods based on pre-trained language models.
We observe that language model-based reasoning methods such as SimKGC [57] and MoCoKGC [65] achieve unusually high results on WN18RR, but perform poorly on other datasets, as shown in Table 7. To better understand this phenomenon, we take these two methods as representative examples for further analysis. We note that WN18RR is derived from WordNet, a large lexical database of English that naturally encodes rich semantic relations between words. Pre-trained language models are well-suited to capturing such general semantic information, which may explain their strong performance on WN18RR. In contrast, FB15k-237 involves more domain-specific relational knowledge, which poses greater challenges for these models, leading to weaker performance (as shown in Table 7).
@@ Appendix D Experimental Details
D.1 Transductive and Inductive Reasoning.
Following the formal definition in [67], transductive reasoning assumes that all test entities appear during training, while inductive reasoning handles completely unseen entities during testing. This difference is fundamental to evaluating model generalization capabilities.
Therefore, following the methodology in [68], we construct our inductive evaluation datasets by ensuring complete separation between training and test entities. This strict partitioning, where test entities are excluded from training, enables a reliable assessment of the model’s generalization capability to unseen knowledge.
D.2 Relation Prediction Task.
The relation prediction task ($h$, ?, $t$) can indeed be transformed to fit our tail completion paradigm through the approach in [69]:
  • Scoring Mechanism: For relation prediction, we fix head ($h$) and tail ($t$) entities, then score all candidate relations. For tail prediction, we fix the head ($h$) and relation ($r$), then score all candidate tails. Both tasks use the same underlying scoring function.
  • Implementation: For relation prediction, we compute a score for each candidate relation and select the one with the highest score as the prediction.
This approach maintains fundamental consistency with tail entity prediction. While the surface-level structures differ, both tasks share the same underlying computational paradigm: evaluating possible completions against fixed components of the triple using a unified scoring mechanism.
D.3 Dataset Statistics
@@ D.4 Hyperparameters Setup
  • Learning rate: $\{10^{-4},\ 5\times 10^{-4},\ 10^{-3},\ 5\times 10^{-3},\ 10^{-2}\}$
  • Weight decay: $\{10^{-5},\ 10^{-4}\}$
  • Hidden dimension: $\{16,\ 32,\ 64,\ 128\}$
@@
In addition, we initialize the value of $\alpha$ randomly within the range (0, 1). Since we use the same random seed for all datasets, the initial value of $\alpha$ is identical across different datasets and is 0.549. And we report the range of $\alpha$ values observed during training across different datasets, as shown in the Table11, Table12 and Table13. The results show that $\alpha$ consistently converges to a stable value during training, with negligible fluctuations afterward (less than 0.001). Across all datasets, $\alpha$ converges reliably as expected. Moreover, $\alpha$ remains below the theoretical threshold $\frac{\sigma_{\max}(\mathcal{M}_{\mathrm{D}})+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}{1+\sigma_{\max}(\mathcal{M}_{\mathrm{O}})}$ as stated in Theorem 1.
Table 11: The range of $\alpha$ values observed during training on FB15k-237.
Datasets FB15k-237v1 FB15k-237v2 FB15k-237v3 FB15k-237v4
Theoretical threshold 2.27 2.44 2.40 2.46
$\alpha$ (Epoch=0) 0.549 0.549 0.549 0.549
$\alpha$ (Epoch=2) 0.538 0.518 0.513 0.507
$\alpha$ (Epoch=4) 0.531 0.501 0.498 0.486
$\alpha$ (Epoch=6) 0.522 0.485 0.477 0.465
$\alpha$ (Epoch=8) 0.514 0.468 0.460 0.450
$\alpha$ (Epoch=10) 0.508 0.457 0.445 0.437
$\alpha$ (Epoch=12) 0.509 0.458 0.445 0.439
$\alpha$ (Epoch=14) 0.511 0.460 0.447 0.440
$\alpha$ (Epoch=16) 0.512 0.461 0.448 0.441
$\alpha$ (Epoch=18) 0.512 0.461 0.449 0.441
$\alpha$ (Epoch=20) 0.512 0.461 0.449 0.441
Table 12: The range of $\alpha$ values observed during training on WN18RR.
Datasets WN18RRv1 WN18RRv2 WN18RRv3 WN18RRv4
Theoretical threshold 2.17 2.00 2.03 2.10
$\alpha$ (Epoch=0) 0.549 0.549 0.549 0.549
$\alpha$ (Epoch=2) 0.539 0.543 0.546 0.543
$\alpha$ (Epoch=4) 0.534 0.531 0.534 0.540
$\alpha$ (Epoch=6) 0.533 0.524 0.536 0.537
$\alpha$ (Epoch=8) 0.532 0.512 0.523 0.536
$\alpha$ (Epoch=10) 0.531 0.504 0.511 0.532
$\alpha$ (Epoch=12) 0.531 0.505 0.511 0.532
$\alpha$ (Epoch=14) 0.533 0.505 0.510 0.533
$\alpha$ (Epoch=16) 0.533 0.505 0.511 0.533
$\alpha$ (Epoch=18) 0.533 0.505 0.511 0.533
$\alpha$ (Epoch=20) 0.533 0.505 0.511 0.533
Table 13: The range of $\alpha$ values observed during training on NELL-995.
Datasets NELL-995v1 NELL-995v2 NELL-995v3 NELL995v4
Theoretical threshold 2.49 2.58 2.54 2.42
$\alpha$ (Epoch=0) 0.549 0.549 0.549 0.549
$\alpha$ (Epoch=2) 0.544 0.536 0.526 0.536
$\alpha$ (Epoch=4) 0.539 0.532 0.496 0.516
$\alpha$ (Epoch=6) 0.523 0.525 0.474 0.500
$\alpha$ (Epoch=8) 0.514 0.517 0.447 0.487
$\alpha$ (Epoch=10) 0.514 0.509 0.424 0.477
$\alpha$ (Epoch=12) 0.514 0.509 0.425 0.475
$\alpha$ (Epoch=14) 0.512 0.510 0.425 0.476
$\alpha$ (Epoch=16) 0.512 0.510 0.426 0.476
$\alpha$ (Epoch=18) 0.512 0.510 0.427 0.476
$\alpha$ (Epoch=20) 0.512 0.510 0.427 0.476
@@
We analyze the impact of the hyperparameter $k$, which denotes the number of entities in the high-score subset. We conduct experiments on the validation sets of all transductive datasets using our model. We set $k$ to {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15}. See Table 15. We ultimately select $k=4$ for all the datasets, as these settings yield relatively high and stable results across three key metrics MRR, Hits@1, and Hits@10 rather than optimizing a single metric in isolation. We also conduct the same experiments on 4 inductive datasets (As shown in Figure 5).
@@
Additionally, We analyze the impact of the hyperparameter, the decision threshold $\Delta$. We run experiments on all transductive datasets with our model. We set $\Delta$ to {0, 0.5, 1, 1.5, 2, 2.5, 3, 3.5, 4, 4.5, 5, 5.5, 6, 6.5, 7, 7.5, 8, 8.5, 9, 9.5, 10}. For FB15k-237 and WN18RR, we set $\Delta$ to 8 and set $\Delta$ to 5 for NELL-995 and YAGO3-10. The results are shown in Table 14.
Table 14: The results on transductive knowledge graph reasoning datasets with different $\Delta$.
Value FB15k-237 WN18RR NELL-995 YAGO3-10
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
0.0 0.411 31.9 58.0 0.581 52.5 68.7 0.553 48.1 66.5 0.600 53.0 72.3
0.5 0.423 33.0 59.2 0.584 53.0 68.9 0.572 50.1 69.3 0.613 54.3 73.0
1.0 0.436 34.1 60.6 0.588 53.4 69.0 0.579 50.8 69.8 0.619 54.8 73.5
1.5 0.443 34.7 61.3 0.589 53.6 69.0 0.585 51.4 70.5 0.623 55.2 74.0
2.0 0.447 35.2 61.7 0.590 53.7 69.1 0.589 51.7 71.1 0.626 55.5 74.4
2.5 0.448 35.3 61.7 0.591 53.8 69.1 0.590 51.8 71.2 0.628 55.7 74.6
3.0 0.450 35.5 61.8 0.591 53.8 69.1 0.592 52.0 71.2 0.630 55.8 74.7
3.5 0.451 35.6 61.9 0.591 53.8 69.2 0.592 52.1 71.2 0.631 56.0 74.8
4.0 0.452 35.7 62.0 0.591 53.9 69.2 0.592 52.1 71.2 0.632 55.1 74.8
4.5 0.452 35.7 62.1 0.591 53.9 69.2 0.592 52.1 71.2 0.632 56.1 74.9
5.0 0.453 35.8 62.2 0.592 53.9 69.2 0.593 52.2 71.2 0.632 56.1 74.9
5.5 0.453 35.8 62.2 0.592 53.9 69.2 0.593 52.2 71.1 0.632 56.1 74.9
6.0 0.454 35.9 62.2 0.592 53.9 69.2 0.593 52.2 71.1 0.632 56.1 74.8
6.5 0.454 35.9 62.3 0.592 53.9 69.2 0.592 52.2 71.1 0.632 56.1 74.8
7.0 0.455 36.0 62.4 0.592 54.0 69.2 0.592 52.2 71.0 0.632 56.1 74.9
7.5 0.456 36.1 62.5 0.593 54.2 69.9 0.592 52.2 71.0 0.632 56.1 74.9
8.0 0.456 36.1 62.8 0.594 54.0 70.0 0.592 52.1 71.0 0.632 56.1 74.9
8.5 0.457 36.1 62.7 0.593 54.0 69.9 0.592 52.1 69.9 0.632 56.0 74.9
9.0 0.456 35.9 62.6 0.593 54.0 69.9 0.592 52.1 69.9 0.632 56.0 74.9
9.5 0.454 35.9 62.6 0.593 54.0 69.2 0.592 52.1 69.9 0.632 56.0 74.9
10.0 0.454 35.9 62.2 0.593 54.1 69.3 0.592 52.1 69.9 0.632 56.0 74.9
@@
D.6 Ranking Protocol.
Following previous work [19], we conducted experiments on DuetGraph using the strictest ranking protocol $(m+n+1)$, where $m$ is the number of entities with higher scores than the correct answer and $n$ is the number of entities that receive the same score as the correct answer. We also conducted experiments using a widely used but more lenient ranking protocols, namely $(m+1)$ adopted in [71, 36, 72, 73], as shown in Table 16 and Table 17.
Table 16: Comparision of different under ranking protocols across four datasets.
Method FB15k-237 WN18RR NELL-995 YAGO3-10
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
KnowFormer ($m+n+1$) 0.430 34.3 60.8 0.579 52.8 68.7 0.566 50.2 67.5 0.615 54.7 73.4
DuetGraph ($m+n+1$) 0.453 36.1 62.4 0.593 54.2 69.9 0.590 52.1 71.2 0.631 56.1 74.8
DuetGraph ($m+1$) 0.456 36.1 62.8 0.594 54.2 70.0 0.593 52.2 71.2 0.632 56.1 74.9
Table 17: Comparison of different ranking protocols across very large knowledge graphs. (Non-deterministic ranking [74, 63, 75] means that when two entities share the same score, their original order is preserved in the ranking. Compared with the $m+n+1$ ranking protocol, this is more lenient.)
Method Wikidata5M Freebase
MRR H@1 H@10 MRR H@1 H@10
AnyBURL (Non-Deterministic Ranking) 0.350 30.9 42.9 0.588 53.6 68.2
KnowFormer ($m+n+1$) 0.332 26.7 46.3 0.684 65.7 73.6
DuetGraph ($m+n+1$) 0.363 32.7 49.5 0.697 69.3 73.8
DuetGraph ($m+1$) 0.363 32.7 49.5 0.699 69.4 73.9
Based on the experimental results, we can draw the following conclusions.
  • First, switching to the strictest ranking protocol has little impact on DuetGraph’s quality. As shown in the tables above, when switching the ranking protocol from $(m+1)$ to $(m+n+1)$, DuetGraph’s quality is virtually unaffected, with at most a 0.3% drop in MRR, a 0.1% drop in Hits@1, and a 0.4% drop in Hits@10. This is because, $n$ is 0 in almost all the cases. As shown in the table above, the number of test triplets impacted when replacing $(m+1)$ with $(m+n+1)$ is less than 1%.
  • Second, DuetGraph achieves SOTA results even under the strictest ranking protocol. As shown above, it surpasses AnyBURL and KnowFormer across all datasets in MRR, Hits@1, and Hits@10. For the other baselines, as shown in Table 2 used the official code from their original papers, all of which employ protocols no stricter than $(m+n+1)$. As shown in [76], increasing the strictness of ranking protocol yields quality that is no higher (and often lower). Even under this tough setting, DuetGraph consistently outperforms all of them, providing strong evidence of its SOTA performance.
Appendix E More Experimental Results
E.1 Model Size and Inference Time.
Despite employing a two-stage reasoning strategy, DuetGraph exhibits significantly superior inference efficiency compared to KnowFormer[19]. This advantage is primarily attributed to our innovative parallel processing architecture and a lightweight coarse-grained model selection mechanism. These components work synergistically, enabling DuetGraph to achieve this higher efficiency at a comparable model scale, as detailed in Table 18.
Table 18: Comparison of Model Size and Inference Time across different datasets.
Datasets FB15k-237 WN18RR NELL-995 YAGO3-10
Metrics Size (M) Time (ms) Size (M) Time (ms) Size (M) Time (ms) Size (M) Time (ms)
DuetGraph
6.5
(Coarse: 0.6, Fine: 5.9)
347.07
1.3
(Coarse: 0.9, Fine: 0.4)
292.49
7.2
(Coarse: 2.3, Fine: 4.9)
261.32
3.4
(Coarse: 2.3, Fine: 1.1)
597.22
KnowFormer[19] 6.1 499.71 0.4 392.40 5.2 360.40 1.0 905.22
As shown in Table 18, the inference efficiency improvement is especially notable on large-scale knowledge graphs, such as YAGO3-10, where our method achieves a 34.2% increase in inference efficiency compared to the SOTA model KnowFormer[19].
E.2 The difference between $\alpha$ and graph attention.
The main differences are reflected in the following three perspectives:
  • Technical Role.: As illustrated in Figure 2, the graph attention mechanism (Step 2) first learns global weights, which subsequently inform the learning of the control parameter $\alpha$ (Step 4). The attention weights serve as intermediate representations that enable $\alpha$ to effectively balance local and global information.
  • Theoretical Advantage.: Introducing $\alpha$ allows for better fusion of the global weights captured by the attention mechanism (Step 2) and the local weights acquired via message passing (Step 3), which helps mitigate over-smoothing and enhances the model’s quality, as shown in Theorem 1.
  • Experimental Study.: Incorporating $\alpha$ achieves better performance compared to using attention mechanism alone, as shown in Table 19.
Table 19: Results comparing the model with and without the $\alpha$ parameter.
Metrics FB15k-237 WN18RR NELL-995 YAGO3-10
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
w/ $\alpha$ 0.456 36.1 62.8 0.594 54.2 70.0 0.593 52.2 71.2 0.632 56.1 74.9
only w/ attention 0.445 35.1 61.2 0.584 54.1 69.0 0.582 51.0 70.3 0.617 54.2 73.7
Appendix F Limitations and Broader Impacts
F.1 Limitations